Find the present value of a series of investment of $1 at the end of each year for 25 years with effective

interest rate of 13.5% p.a.


1
Expert's answer
2021-05-07T11:28:02-0400

"=P[\frac{1 - (1 +r) ^{-n}} {r}]"


"=\$1[\frac{1 - (1 +0.00135) ^{-25}} {0.00135}]"


"=\$1 [\frac{1 - (1.00135) ^{-25}} {0.00135 }]"


"=\$1 [\frac{1 - (1 - 0.967)} {0.00135 }]"


"=\$1 (24.44)"


Present Value ="\$24.44"